1)3x² + 5x – 12 = 0
2)2x² + 2x – 1 = 0

Answer by checkley75(3666)

3X^2+5X-12=0
(3X-4)(X+3)=0
3X-4=0
3X=4
X=3/4 ANSWER.
X+3=0
X=-3 ANSWER.
--------------------------------------------
2X^2+2X-1=0
USING THE QUADRATIC EQUATION WE GET
X=(-B+-SQRT[B^2-4*A*C])/2A
X=(-2+-SQRT[4-4*2*-1])/2*2
X=(-2+-SQRT[4+8])/4
X=(-2+-SQRT12)/4
X=(-2+-3.464)/4
X=(-2+3.464)/4
X=1.464/4
X=.366 ANSWER
X=(-2-3.464)/4
X=-5.464/4
X=-1.366 ANSWER